# Remote Cooling of Spin-ensembles through a Spin-mechanical Hybrid Interface

### v2

@@ Abstract
We present a protocol for the ground-state cooling of a tripartite hybrid quantum system, in which a macroscopic oscillator acts as a mediator between a single probe spin and a remote spin ensemble. In the presence of weak dispersive coupling between the spins and the oscillator, cooling of the oscillator and the ensemble spins can be achieved by exploiting the feedback from frequent measurements of the single probe spin. We explore the parameter regimes necessary to cool the ensemble, the oscillator, or both to their thermal ground states. This novel cooling protocol shows that, even with only weak dispersive coupling, energy transfer-like effects can be obtained by simply manipulating the probe spin. These results not only contribute to the development of a practical solution for cooling/polarizing large spin ensembles, but also provide a relatively simple means of tuning the dynamics of a hybrid system. The proposed cooling protocol thus has broader implications for advancing various quantum technology applications, such as macroscopic quantum state generation and remote sensing.
@@ I Introduction
Translating the potential of quantum technology into real-world applications remains a significant challenge due to the inherent limitations of current experimental platforms [1, 2]. One promising avenue is to explore the hybrid system approach, which promises to fully exploit the unique capabilities of different systems [3, 4, 5, 6]. In particular, high-fidelity entanglement between different systems can be established by exploiting their common coupling to a transducer, thereby generating indirect interactions [7, 6]. Transducers in the forms of nanomechanical oscillators (NMOs) [8, 9], magnons [10, 11], superconducting qubits [12, 13], superconducting resonators [14], optical cavities [15], and optomechanical cavities [16] have been considered to mediate spin–spin, spin-photon, and photon-phonon interactions.
However, it has been mainly proposed to use transducers to implement entanglement gates between remote qubits [16, 9, 10], and few transducer protocols focus on heat extraction from quantum systems to achieve ground state cooling or state preparation. The development of efficient polarization methods is particularly critical for spin ensembles, as conventional techniques often rely on resonant energy exchange methods, such as the Hartman-Hahn transfer [17, 18]. Polarization based on energy transfer typically exploits the interaction between a single spin and all the spins of the ensemble. This single spin can be polarized repeatedly, thus realizing a continuous removal of entropy until the ensemble is finally polarized. The challenge is that if the ensemble spins have similar g-factors, the risk of them being trapped in dark states due to Hamiltonian symmetries is high, preventing complete polarization of the ensemble [18, 19].
@@
The oscillator here serves as a simplified conceptual model for a common bus (transducer) connecting two distant quantum systems that are not directly or weakly coupled. One can consider a general scenario where the bus is a large spin network that can be effectively approximated as a bosonic system via the Holstein-Primakoff transformation, such as the spin wave excitations (magnons) [22, 11]. This is relevant when a single spin, such as nitrogen-vacancy (NV) centers in diamond [23, 24], is used to polarize or probe distant target spins beyond the distance criteria set by the sensing volume [25]. The target may be distant carbon nuclear spins within the diamond substrate [26, 27], or spins attached to an external protein/molecule [28]. In such cases, probe-target interactions must be mediated by a transducer, such as the spin network schematically shown in the inset of Fig. 1(a) [26, 27, 28, 10]. Despite tuning both the network-to-probe and network-to-target interactions to resonance, the significant inhomogeneities typically present within a spin network make it difficult to become polarized, rendering resonant probe-target energy transfer ineffective [28, 26, 27]. Here, we instantiate such a scenario by considering the oscillator as the fundamental mode of an NMO, specifically the vibration of a clamped cantilever, as schematically shown in Fig. 1(a) [20, 21].
The paper is organized as follows: In Sec. II, we outline the Hamiltonian of the considered hybrid system and discuss possible experimental implementations of the proposed protocol. In Sec. A, we describe the dynamics of the hybrid system and analyze how periodic inversion pulses on the probe spin can significantly modify the dynamics of the ensemble and the oscillator. In Sec. IV, we detail how repeated probe projections can bring the oscillator and the ensemble into their ground states. In particular, we study the parameters that allow cooling of the spin ensemble, the oscillator, or both. Additionally, we discuss how it is possible to prepare the oscillator into a complex quantum state using the same probe control sequence as the cooling protocol. Finally, Sec. V provides a summary and discusses possible directions for future research.
@@ II System Hamiltonian and experimental realization
where $S_{i,z}$ signifies the z-component of the spin-1/2 operator for the ith spin, oscillating at the Larmor frequency $\omega_{i}$. The oscillator’s annihilation and creation operators are denoted by $a$ and $a^{\dagger}$, with a frequency $\omega$.
@@
Figure 1: Schematic representation of the experimental setup and control strategy. (a) Shown is a clamped nano/micromechanical cantilever carrying a single defect spin, which in this study is exemplified by an NV electronic spin. The interaction between the NV center spin and the cantilever motion is mediated by strain, while single qubit gates are achieved by microwave (MW) pulses. Initialization and readout processes are facilitated by optical pulses, with a photon detector used for readout. A magnetic tip at the edge of the cantilever creates a gradient field that allows interaction with a nearby spin ensemble, coupling it to the mechanical motion of the cantilever. In the inset, we schematically represent a network of interacting spins as a quantum bus. When the number of spins in the network is large, it can be modeled as a bosonic system (e.g., magnons) by the Holstein-Primakoff transformation. Note that it is simplified here as the fundamental mode of the cantilever’s vibration. (b) The control sequence of the probe spin to perform the cooling protocol. First, the probe spin is reset to the |$\ket{0}$ state by optical pulses, then a $\pi/2$ MW pulse brings it to a state of equal superposition, which is followed by $n$ inversion MW pulses applied periodically with a carefully chosen time interval $\tau$. Another $\pi/2$ MW pulse is applied before the probe is optically read into the computational base. The above process is repeated $M$ times, bringing the oscillator and the spin ensemble progressively closer to their thermal ground states. Note that the probe to be read must be post-selected in |$\ket{0}$ each time, otherwise the whole process starts over.
@@
This hybrid system can be realized as shown in Fig. 1(a) by a unilaterally clamped cantilever, embedding a single NV center positioned at the clamping point such that it experiences maximum strain and thus higher coupling strength. On the free side of the cantilever, we attach a magnetic tip to generate a strong field gradient that couples to a nearby spin ensemble on an external substrate [9, 6]. In this way, we exploit both possibilities to achieve spin-mechanical coupling [7, 29, 30] by local (strain) and non-local (magnetic-tip) coupling of spins to mechanical modes within a single setup. This allows the distinct manipulation of quantum systems that are otherwise not controllable.
The NV electronic spin is spin-1, which can be operated as a qubit with its $m_{s}=0,-1$ sublevels. This requires an external magnetic field aligned with its symmetry axis to separate the $m_{s}=\pm 1$ levels, which is typically around 400 Gauss [32]. Such an arrangement negates the need to consider transverse spin-strain coupling due to its direct interaction between the $m_{s}=\pm 1$ levels [7, 29, 30], allowing us to safely ignore the $m_{s}=+1$ sublevel in this work. What remains is the longitudinal strain coupling acting on the qubit levels, with a number that can be up to 10 kHz for NV centers [36]. This number can be even higher for other solid-state defect centers [37].
In contrast, the spin ensemble to be cooled can be chosen more flexibly, since precise control of these spins is not a requirement for the current study. Potential targets include distant spins on other coupled cantilevers, spins at the clamped edge that are strain-coupled to the mechanical mode, or surface electron spins that are not well controlled. For example, if the ensemble spins are dark spins on the surface of the NMO [38] or other ferromagnetic particles [39], optical manipulation over them can be difficult, highlighting the need for other indirect polarization techniques. Note that this ensemble-oscillator interaction generated by the attached magnetic tip can be (1-2 orders of magnitude) stronger than the spin-strain coupling between the probe and the oscillator [40].
@@ A Original system dynamics
In the following, we will consider the system in the rotating frame of precession frequencies of the oscillator and the spins. The rotating-frame Hamiltonian reads
@@
Since the ensemble spins form a finite-dimensional Hilbert space, it is convenient to write the system evolution operators in the form [42, 43]:
@@
$h_{\scriptscriptstyle k,\pm}(t)=\frac{g(k-N/2)\pm g_{0}}{2}\left(\hat{a}e^{-i\omega t}+\hat{a}^{\dagger}e^{i\omega t}\right);$ (4)
@@
$\begin{split}\mathcal{D}_{\scriptscriptstyle k,\pm}(t&)=\exp\left(\pm i\int_{0}^{t}\mathrm{d}t^{\prime}h_{\scriptscriptstyle k,\pm}(t^{\prime})\right)\\ &\qquad\times\exp\left(\frac{1}{2}\int_{0}^{t}\mathrm{d}t^{\prime}\int_{0}^{t^{\prime}}\mathrm{d}t^{\prime\prime}\left[h_{\scriptscriptstyle k,\pm}(t^{\prime}),h_{\scriptscriptstyle k,\pm}(t^{\prime\prime})\right]\right)\\ &=\exp\left[\frac{g(k-N/2)\pm g_{0}}{2\omega}\left(\alpha(t)a^{\dagger}-\alpha^{*}(t)a\right)\right]\\ &\qquad\quad\times\exp\big{(}it\theta_{\scriptscriptstyle k}\big{)}.\end{split}$ (6)
@@
B Inversion pulses on probe
In addition to the natural dynamics induced by the Hamiltonian given in Eq. (1), we also apply a periodic sequence of $\pi$ pulses to the probe spin. These pulses significantly alter the dynamics of the system by imposing a time-periodic function:
$f(t)=\begin{cases}+1&\text{if }2j\tau<t<(2j+1)\tau\\ -1&\text{if }(2j+1)\tau<t<(2j+2)\tau\end{cases}$ (8)
where $j=0,1,2,\cdots,n$, and $\tau$ denotes the interval between successive pulses. By replacing $g_{0}$ with $f(t)g_{0}$ in $h_{\scriptscriptstyle k,\pm}(t)$, the conditional evolution operators in Eq. (6) are now written as
@@ B Inversion pulses on probe
To achieve some tunability of the system dynamics that facilitates our proposed cooling protocol, we introduce the pulse detuning $\epsilon$ as a control parameter:
@@
For higher values of pulse detuning $\epsilon$, making the above equation closer to a delta function would require an even larger pulse number $n$.
Figure 2: (a) The quantum circuit progressively projects the oscillator ensemble system by projection of the probe spin. The pulse sequence on the physical level is shown in fig. 1(b). In each repetition, the probe spin is first prepared in the equal superposition state $\ket{+}$ by the optical reset and a microwave Hadamard gate. Then repeated inversion pulses are applied, resulting in a controlled oscillator displacement $\sum_{k}\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)\otimes\mathrm{I}_{k}$. Here the subscript $\pm$ depends on the spin state of the probe and $k$ on the number of upward-pointing ensemble spins. The procedure ends with a Hadamard gate and a measurement of the probe spin, with the result set to 0. Such a repetition effectively implements the projector given in Eq. (14). (b) An example of how an ensemble of $N=14$ spins is polarized after $M=100$ probe projections. The spins are initially in a completely mixed state, which is projected to the state with all spins pointing up with a probability of 95.82%. Note that achieving such a high degree of polarization requires a carefully chosen pulse detuning, which in this case is $\epsilon=1.16$ kHz. Further simulation details are summarized in Appendix A.
C Tuning dynamics through probe projections
Although we focus specifically on the weak dispersive coupling between each spin and the oscillator, we deliberately ignore any direct interactions between the spins, making energy transfer between the probe and the ensemble unlikely. Nevertheless, by carefully choosing the pulse detuning $\epsilon$ and the number of pulses $n$, energy transfer-like changes in the dynamics of the spin ensemble can be induced, as we will discuss below.
The dynamics modification is facilitated by a cyclic process, involving the projective measurement and reinitialization of the probe spin, as schematically shown in Fig. 1(a). First, the probe is prepared in the state $\ket{+}$. Then the system undergoes an evolution for a time $t=n\tau$ under repeated inversion pulses, as described in Eq (3). Finally, the probe spin is measured in the computational basis after an additional $\pi/2$ pulse is applied to it.
The process described above is represented by the circuit in Fig. 2(a), effectively implementing oscillator displacements conditioning on the probe state. The whole circuit is then equivalent to the implementation of a projector on the oscillator-ensemble system, which is mathematically formulated as:
$\mathcal{P}_{\scriptscriptstyle N,\pm}=\sum_{\scriptscriptstyle k=0}^{\scriptscriptstyle N}V_{k,\pm}\otimes\mathrm{I}_{k},\quad V_{k,\pm}=\frac{1}{2}\left(\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,+}(t)\pm\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,-}(t)\right).$ (14)
The operator $V_{k,\pm}$ is a projector on the oscillator, where $k$ is the number of ensemble spins pointing up, and the subscript $\pm$ depends on the measurement result of the probe spin.
Note that such repeated projections of the probe is central to the oscillator cooling protocol in our previous work [43], which is now extended to cooling/polarizing a spin ensemble. Here, we start with $N$ ensemble spins initially in a fully mixed state, and the oscillator initially in its thermal state. These initial states are written as
[…]
To cool the ensemble-oscillator subsystem, we repeatedly implement the circuit shown in Fig. 2(a). It is important to emphasize that, each time, only a probe measurement result of 0 is accepted, otherwise the whole process is restarted. This process effectively implements the projector $\mathcal{P}_{\scriptscriptstyle N,+}$ given in Eq. (14) many times. The probability that the ensemble is projected into the $\mathrm{I}_{k}$ sector after $M$ repetitions is calculated as

### v3

@@ Abstract
We present a protocol for the ground-state cooling of a tripartite hybrid quantum system, in which a macroscopic oscillator acts as a mediator between a single probe spin and a remote spin ensemble. In the presence of weak dispersive coupling between the spins and the oscillator, cooling of the oscillator and the ensemble spins can be achieved by exploiting the feedback from frequent measurements of the single probe spin. We explore the parameter regimes necessary to cool the ensemble, the oscillator, or both to their thermal ground states. This novel cooling protocol shows that, even with only weak dispersive coupling, energy transfer-like effects can be obtained by simply manipulating the probe spin. These results not only contribute to the development of a practical solution for cooling/polarizing large spin ensembles, but also provide a relatively simple means of tuning the dynamics of a hybrid system. The proposed protocol thus has broader implications for advancing various quantum technology applications, such as macroscopic quantum state generation and remote sensing.
@@ I Introduction
Translating the potential of quantum technology into real-world applications remains a significant challenge due to the inherent limitations of current experimental platforms [1, 2]. One promising avenue is to explore the hybrid system approach, which promises to fully exploit the unique capabilities of different systems [3, 4, 5, 6]. In particular, high-fidelity entanglement between different systems can be established by exploiting their common coupling to a transducer, thereby generating indirect interactions [7, 6].
@@
The oscillator here serves as a simplified conceptual model for a common bus (transducer) connecting two distant quantum systems that are not directly or weakly coupled. One can consider a general scenario where the bus is a large spin network, which can be effectively approximated as a bosonic system via the Holstein-Primakoff transformation, such as the spin wave excitations (magnons) [10, 11]. This scenario is particularly relevant when a single nitogen-vacancy (NV) electronic spin in diamond is used to polarize or probe distant target spins beyond the distance criteria set by the sensing volume [12]. The target may be distant carbon nuclear spins within the diamond substrate [13, 14, 15], or spins attached to an external protein/molecule [16]. In such cases, probe-target interactions must be mediated by a transducer, such as the spin network schematically shown in the inset of Fig. 1(a) [14, 15, 16, 17]. Despite tuning both the network-to-probe and network-to-target interactions to resonance, the significant inhomogeneities typically present within a spin network make it difficult to become polarized, rendering resonant probe-target energy transfer ineffective [16, 14, 15]. Here, we instantiate such a scenario by considering the oscillator as the fundamental mode of an NMO, specifically the vibration of a clamped cantilever, as schematically shown in Fig. 1(a) [8, 9].
The paper is organized as follows: In Sec. II, we outline the Hamiltonian of the considered hybrid system and discuss possible experimental implementations of the proposed protocol. In Sec. III, we describe the dynamics of the hybrid system and analyze how periodic inversion pulses on the probe spin can significantly modify it. In Sec. IV, we detail how repeated probe projections can bring the oscillator and the ensemble into their ground states. In particular, we demonstrate the effectiveness of our protocol by investigating the parameters that allow cooling of the spin ensemble, the oscillator, or both. Additionally, we explore how the same probe control sequence used for cooling can be adapted to prepare the oscillator in complex quantum states. Finally, Sec. V summarizes our findings and outlines potential avenues for future research.
@@ II System Hamiltonian and experimental realization
where $N$ is the total number of spins in the ensemble; $S_{i,z}$ signifies the z-component of the spin-1/2 operator for the ith spin, oscillating at its Larmor frequency $\omega_{i}$. The oscillator’s annihilation and creation operators are denoted by $a$ and $a^{\dagger}$, with a frequency $\omega$.
As we will demonstrate later, the assumption of only a single bosonic mode is valid in this work. This validity stems from our manipulation of the probe spin through periodic inversion pulses, which effectively couples the probe spin exclusively to the oscillator’s ground state mode while decoupling it from all others. Moreover, our focus on a small ensemble of only a few spins allows for exact diagonalization of the system Hamiltonian. This approach stands in contrast to typical acousto-magnonic studies, where a bosonic description of much larger spin ensembles is necessary [18]. Such bosonic approximations are, however, inapplicable in our few-spin limit.
@@
Figure 1: Schematic representation of the experimental setup and control strategy. (a) Shown is a clamped nano/micromechanical cantilever carrying a single defect spin, which in this study is exemplified by an NV electronic spin. The interaction between the NV center spin and the cantilever motion is mediated by strain, while single qubit gates are achieved by microwave (MW) pulses. Initialization and readout processes are facilitated by optical pulses, with a photon detector used for readout. A magnetic tip at the edge of the cantilever creates a gradient field that allows interaction with a nearby spin ensemble, coupling it to the mechanical motion of the cantilever. In the inset, we schematically represent a network of interacting spins as a quantum bus. When the number of spins in the network is large, it can be modeled as a bosonic system (e.g., magnons) by the Holstein-Primakoff transformation. Note that it is simplified here as the fundamental mode of the cantilever’s vibration. (b) The control sequence of the probe spin to perform the cooling protocol. First, the probe spin is reset to the |$\ket{0}$ state by optical pulses, then a $\pi/2$ MW pulse brings it to a state of equal superposition, which is followed by $n$ inversion MW pulses applied periodically with a carefully chosen time interval $\tau$. To ensure that the effective probe-oscillator coupling strength under pulses is comparable to the oscillator frequency $\omega$, enabling the oscillator to be driven, $n$ should be on the order of $\omega/g_{0}$. Another $\pi/2$ MW pulse is applied before the probe is optically read into the computational base. The above process is repeated $M$ times, bringing the oscillator and the spin ensemble progressively closer to their thermal ground states. Note that the probe to be read must be post-selected in |$\ket{0}$ each time, otherwise the whole process starts over.
@@
This hybrid system can be realized as shown in Fig. 1(a) by a unilaterally clamped cantilever, embedding a single NV center positioned at the clamping point such that it experiences maximum strain and thus higher coupling strength. On the free side of the cantilever, we attach a magnetic tip to generate a strong field gradient that couples to a nearby spin ensemble on an external substrate [19, 6]. In this way, we exploit both possibilities to achieve spin-mechanical coupling [7, 20, 21] by local (strain) and non-local (magnetic-tip) coupling of spins to mechanical modes within a single setup. This allows the distinct manipulation of quantum systems that are otherwise not controllable. More details about this possible experimental setup can be found in the Appendix. A.
The NV-cantilever setup we’ve outlined serves as a concrete example of the broader, abstract challenge of harnessing noisy environments. We selected this particular system for two primary reasons. First, achieving polarization is a crucial objective in quantum technology applications using solid-state spins such as NV centers, especially in scenarios where resonant energy transfer is improbable [22, 23, 24]. Second, and perhaps more significantly, this setup has already been successfully implemented experimentally [25, 21], providing a solid foundation for further investigation of using noisy environment as resources.
@@ III System dynamics and its tuning
In this section, we provide analytical derivations of the system dynamics and illustrate how periodic inversion pulses on the probe spin can significantly affect the behavior of the whole system when certain resonance conditions are roughly met. These derivations are intended to provide insight into the mechanisms of our cooling protocol shown schematically in Fig. 1(b).
@@ A Original system dynamics
In the following, we will consider the system in the rotating frame with respect to the precession frequencies of both the oscillator and the spins. This transformation is applied to an arbitrary quantum state |$\ket{\psi}$ as follows: $\exp{(-i\omega a^{\dagger}a)}\exp{(-i\sum_{k}\omega_{k}S_{k,z})}\ket{\psi}$ [26]. Then the rotating-frame Hamiltonian reads
@@
This Hamiltonian allows us to study the effect of our cooling protocol on each spin configuration with a given eigenvalue of the collective spin operator $S^{z}=\sum_{k=1}^{N}S_{k,z}$, independently. The system evolution operators are then conveniently decomposed into the following form [27, 28]:
@@
$h_{\scriptscriptstyle k,\pm}(t)=\frac{g(2k-N)\pm g_{0}}{2}\left(\hat{a}e^{-i\omega t}+\hat{a}^{\dagger}e^{i\omega t}\right);$ (4)
@@
$\begin{split}\mathcal{D}_{\scriptscriptstyle k,\pm}(t&)=\exp\left(\pm i\int_{0}^{t}\mathrm{d}t^{\prime}h_{\scriptscriptstyle k,\pm}(t^{\prime})\right)\\ &\qquad\times\exp\left(\frac{1}{2}\int_{0}^{t}\mathrm{d}t^{\prime}\int_{0}^{t^{\prime}}\mathrm{d}t^{\prime\prime}\left[h_{\scriptscriptstyle k,\pm}(t^{\prime}),h_{\scriptscriptstyle k,\pm}(t^{\prime\prime})\right]\right)\\ &=\exp\left[\frac{g(2k-N)\pm g_{0}}{2\omega}\left(\alpha(t)a^{\dagger}-\alpha^{*}(t)a\right)\right]\\ &\qquad\quad\times\exp\big{(}it\theta_{\scriptscriptstyle k}\big{)}.\end{split}$ (6)
@@
The analysis can be significantly simplified by considering each spin component with a fixed value of $k$ separately. The initial ensemble state is diagonal in the basis of the collective spin operator $S^{z}$, and this diagonal nature is preserved throughout the complete system evolution. Consequently, each $k$-dependent angle $\theta_{k}$ becomes a global phase for the corresponding spin component with a particular eigenvalue of $S^{z}$, allowing us to disregard it in our analysis. It is also important to note that the oscillator-induced spin-spin interaction, plays no role in the cooling process, as it also takes the form $\propto\sum_{i,j}{S_{i}^{z}S_{j}^{z}}$ [29].
B Dynamical Decoupling pulses on probe
The original system dynamics are unlikely to exhibit energy transfer-like effects due to the weak and off-resonant spin-oscillator couplings. To introduce some tunability into these dynamics, we consider applying a periodic sequence of $\pi$ pulses to the probe spin. Our cooling method has some similarity to how high-fidelity entangling gates can be realized between an NV electronic spin and a nuclear spin, which is also achieved by periodic inversion pulses on the electronic spin [30].
Similar to the nulear-spin polarization schemes [26, 31]), our cooling protocol relies on significantly altering the system dynamics by fine-tuning the pulses on the probe spin to match the oscillator freqeuncy. This allows the effects of weak spin-oscillator couplings to accumulate over time, thereby enabling effective cooling of the spin ensemble and the oscillator. These pulses alter the dynamics of the system by replacing $g_{0}$ with $f(t)g_{0}$ in Eq. (4). The function $f(t)$ is time-periodic and expressed as
$f(t)=\begin{cases}+1&\text{if }2j\tau<t<(2j+1)\tau\\ -1&\text{if }(2j+1)\tau<t<(2j+2)\tau\end{cases},$ (8)
where $j=0,1,2,\cdots,n$ and $\tau$ denotes the interval between successive pulses. The conditional evolution operators in Eq. (6) are now written as
@@ B Dynamical Decoupling pulses on probe
The average displacement magnitude caused by the ensemble spins can be approximated as $\alpha_{E}(k)\approx g(k-N/2)/2\omega$, which is notably small compared to the probe-induced displacement. The latter, given by $\alpha_{P}\approx ng_{0}/\omega$, scales with the number of pulses $n$ when the resonance condition is met, as shown in Eq. (11). Consequently, in a free induction decay experiment of the probe spin, where no inversion pulses are applied, the ensemble-induced effects on the oscillator remain undetectable. However, in our protocol, each probe projection modifies the oscillator state, conditioning it on these minute ensemble-induced displacements. Although initially negligible, these effects accumulate over the repeated probe projections, ultimately exerting a significant influence on the system dynamics.
Figure 2: (a) The quantum circuit progressively projects the oscillator ensemble system by projection of the probe spin. The pulse sequence on the physical level is shown in fig. 1(b). In each repetition, the probe spin is first prepared in the equal superposition state $\ket{+}$ by the optical reset and a microwave Hadamard gate. Then repeated inversion pulses are applied, resulting in a controlled oscillator displacement $\sum_{k}\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)\otimes\mathrm{I}_{k}$. Here the subscript $\pm$ depends on the spin state of the probe and $k$ on the number of upward-pointing ensemble spins. The procedure ends with a Hadamard gate and a measurement of the probe spin, with the result set to 0. Such a repetition effectively implements the projector given in Eq. (12). (b) An example of how an ensemble of $N=14$ spins is polarized after $M=100$ probe projections. The spins are initially in a completely mixed state, which is projected to the state with all spins pointing up with a probability of 95.82%. Note that achieving such a high degree of polarization requires a carefully chosen pulse detuning, which in this case is $\epsilon=1.16$ kHz. Further simulation details are summarized in Appendix C.
C Tuning dynamics through probe projections
In this section, we show how the system dynamics can be modulated by applying inversion pulses to the probe spin. The dynamics modification is facilitated by a cyclic process involving projective measurement and reinitialization of the probe spin, as schematically shown in Fig. 1(a).
First, the probe is prepared in the state $\ket{+}$. Then the system undergoes an evolution for a time $t=n\tau$ under repeated inversion pulses as described in Eq (3). Finally, the probe spin is measured in the computational basis after an additional $\pi/2$ pulse is applied to it. This process is represented by the circuit in Fig. 2(a), which effectively implements oscillator displacement conditioning on the probe spin state.
The whole circuit is equivalent to the implementation of projection operators on the oscillator that depend on the probe spin as well as the ensemble spin states. These conditional oscillator projection operators $V_{k,\pm}$ are mathematically formulated as:
$\mathcal{P}_{\scriptscriptstyle N,\pm}=\sum_{\scriptscriptstyle k=0}^{\scriptscriptstyle N}V_{k,\pm}\otimes\mathrm{I}_{k},\quad V_{k,\pm}=\frac{1}{2}\left(\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,+}(t)\pm\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,-}(t)\right),$ (12)
where $k$ is the number of ensemble spins pointing up, and the subscript $\pm$ depends on the measurement result of the probe spin. Such repeated projections of the probe are central to the oscillator cooling protocol in our previous work [28], which is now extended to cooling/polarizing a spin ensemble.
We start with an ensemble of $N$ spins initially in a fully mixed state and the oscillator initially in its thermal state. These initial states are written as
[…]
To cool the ensemble-oscillator subsystem, we repeatedly implement the circuit shown in Fig. 2(a) by post-selecting the probe measurement outcome to be 0 each time. This process effectively implements the projector $\mathcal{P}_{\scriptscriptstyle N,+}$ given in Eq. (12) many times. The probability that the ensemble is projected into the $\mathrm{I}_{k}$ sector after $M$ repetitions is calculated as
D Pulse detuning
To achieve tunability of the system dynamics described in Eq. (14) and facilitate our proposed cooling protocol, we introduce the pulse detuning $\epsilon$ as a control parameter:
@@ D Pulse detuning
For higher values of pulse detuning $\epsilon$, making the above equation closer to the delta function given in Eq. (11) would require an even larger pulse number $n$.